by TEI TEI

01/03/06 – TTC received final paperwork on this project. TE will process it and turn it into Tony for closeout on this project.

12/02/05 – TTC contacted Process again for final paperwork.

11/28/05 – TTC was at the CH today and discussed with Mike Bartlett final job status and any punchlist items. MB did not report any further punchlist items and TTC did not record any incomplete from original walk through that asked for the insulation to be complete and the programming and scaling of the steam flow from the new meter to be re-adjusted. These issues are complete and therefore no further punchlist items are required at this time. TTC approved allowance usages to repair a defective gasket on an isolation valve during the project for $1000 which included some miscellaneous other travel charges and such to get parts for this replacement and a start/stop scenario while ThermalTech gathered information to allow replacement to take place. TTC will contact process for final payment.

11/08/05 – Mark Neary reprogrammed the Teletrol system, Mike Bartlett confirmed the flow readings were correct. Insulation on the old steam flow meter spool piece is not complete yet but everything seems to be working fine.

11/07/05 – Process sent in pencil copy payapp. TTC contacted Mark Neary for online programming of the Teletrol trying to save paying a trip from Columbus for work.

11/01/05 – Project is complete aside from reprogramming the Teletrol system to the new range of inputs from the low flow meter. TTC is working with Mark Neary to get him to dial in and change the expected input to match the actual flow. Mike Bartlett reported today that the partial flow was around 1900 lbs/hr.

10/24/05 – TTC emailed Mike & Clem to get project update.

10/21/05 – Mike reported that the main steam flow was back up and running for the JC. The flow meter eccentric reducers were being welded in. One was already installed.

10/18/05 – Process is all set to begin and finish work this week on the project. The steam plant is down and cooling off. TTC will stop at the site on Wed morning to coordinate and then go to a prebid on a different project. Process will attempt to do the work on the existing steam meter first and get the plant back up and running on Wednesday.

10/14/05 – Personnel issues are preventing Process from starting but Clem reported that they can start on Wed 10/19 at the CH.

10/04/05 – We were trying to be done with the Steam Flow meter the first week of October. TTC has contacted Process about availability of parts and getting this work completed.